OpenStudy (anonymous):

solve and check 2x^2+2x=4

Parth (parthkohli):

I'd probably start by dividing both sides by 2 for reducing the work. :)

Parth (parthkohli):

C'mon, give it a try.

OpenStudy (anonymous):

2x^2/2 +2x=4/2

Parth (parthkohli):

You should get this:\[x^2 + x = 2 \implies x^2 + x - 2 = 0 \]There's a quadratic equation.

OpenStudy (anonymous):

what do I have to do next?

OpenStudy (anonymous):

the answer is 2, 0